# CMAR-Net: Accurate Cross-Modal 3D SAR Reconstruction of Vehicle Targets with Sparse Multi-Baseline Data

### v3

@@ 3.2.4 Feature space interpolation
3.3 Results for real-world data
3.3.1 Real-world dataset
The real-world data used in this study is sourced from the GOTCHA Circular SAR dataset, collected and released by the United States Air Force Laboratory[39]. This dataset captures a scene with various civilian vehicles using a radar with the center frequency of 9.6 GHz and the bandwidth of 640 MHz. The data is acquired in circular SAR mode through 8 passes at different altitudes, with elevation angles of [45.66°, 44.01°, 43.92°, 44.18°, 44.14°, 43.53°, 43.01°, 43.06°]. The combination of circular observation aperture and altitude diversity enabled 3D imaging of the entire scene. We select data from two SUVs and two sedans. Figure 16 illustrates the scene, radar passes, and vehicle photos.
@@ 3.3.1 Real-world dataset
3.3.2 Imaging results
Figure 17 to 21 present the imaging results on real-world data with varying number of sub-apertures. The BP method produces results heavily affected by noise and side lobes, which can hardly be recognized. The CS method alleviates these issues, but the reconstructions still appear as chaotic point clouds, leaving vehicle features unrecognizable, especially when using highly sparse data. For example, when the number of sub-apertures is 4, traditional methods completely fail, producing totally disordered point clusters without any discernible or meaningful vehicle features.
In contrast, radar image-supervised DL methods can reconstruct vaguely discernible vehicle contours. However, as the number of apertures decreases, these contours deteriorates. For example, when the number of sub-apertures is reduced to 4, SACNet and UNet3D show significant omissions in reconstructing the roof of Jeep. After integrating cross-modal supervision, SACNet and UNet3D exhibit weak generalization when processing real-world data, struggling to accurately reconstruct the vehicle’s geometry. Particularly, when the number of sub-apertures is 4, SACNet fails completely and cannot output any visible features.
Figure 17: Imaging results of different methods using real-world data. Imaging using 12 sub-apertures
Figure 18: Imaging results of different methods using real-world data. Imaging using 10 sub-apertures
Figure 19: Imaging results of different methods using real-world data. Imaging using 8 sub-apertures
Figure 20: Imaging results of different methods using real-world data. Imaging using 6 sub-apertures
Figure 21: Imaging results of different methods using real-world data. Imaging using 4 sub-apertures
@@ 3.3.2 Imaging results
Methods Number of sub-aperture
@@
Methods Number of sub-aperture
@@
Figure 22: Image quality metrics-number of sub-aperture curves for different methods on real-world data.
Tables 6-7 and Figure 22 show the evaluation metrics. Compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of $>$100% in IoU, 89.36% in CE, 75.83% in PSNR and 47.85% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 86.42% in IoU, 67.99% in CE, 38.81% in PSNR and 10.75% in SSIM.
3.3.3 Ablation study
The 7th and 8th sets of data in the Figure 17-21 and table 6-7 show the imaging results and evaluation metrics after ablating the PRP unit and data augmentation strategy. Without the PRP unit, the vehicle images become blurry and suffer greater pixel loss as the number of apertures decreases. And in the absence of data augmentation, the network consistently reconstructs all vehicle type as sedan type, revealing severe overfitting. The PRP unit design improves SSIM, PSNR, IoU, and CE by 8.25%, 21.25%, 24.15%, and 42.82%, respectively. Data augmentation contributes improvements of 8.65%, 19.92%, 42.94%, and 46.75%. These results clearly demonstrate the effectiveness of both the PRP unit and data augmentation strategy.
4 Conclusion
In this paper, we propose a DL-based multi-baseline SAR 3D reconstruction method that utilizes only sparse aperture data. We introduce the concept of cross-modal learning and design a deep neural network tailored for SAR 3D reconstruction—CMAR-Net. By leveraging differentiable rendering, CMAR-Net can be effectively trained, addressing the key challenge of supervising 3D imaging network with 2D optical images. Extensive experiments on both simulated and real-world datasets show that CMAR-Net significantly outperforms all existing baseline methods, including traditional imaging algorithms and DL-based approaches, in terms of image quality and generalization. Moreover, the existing DL-based SAR reconstruction methods, adopting our proposed cross-modal learning paradigm, experience a significant performance improvement, which highlights the superiority and versatility. Last but not least, our method eliminates the need for time-consuming preprocessing of full-aperture data. Instead, it only needs 2D images obtained through computer rendering or camera capture. This significantly reduces the difficulty of dataset construction, greatly enhancing the practicality of SAR 3D reconstruction technology.
Our approach offers new insights for 3D SAR researchers and is expected to comprehensively improve the research landscape of SAR 3D imaging. The technique has the potential to be applied to fields such as aerial remote sensing, terrain mapping, and obstacle detection. In the future, researchers can expand the scope to 3D reconstruction of other scenes (e.g., buildings, streets), or explore fusion with other heterogeneous data (e.g., infrared data), and so on.

### v4

@@ III-B4 Feature space interpolation
III-C Results for real-world data
III-C1 Real-world dataset
The real-world data used in this study is sourced from the GOTCHA Circular SAR dataset, collected and released by the United States Air Force Laboratory[43]. This dataset captures a scene with various civilian vehicles using a radar with the center frequency of 9.6 GHz and the bandwidth of 640 MHz. The data is acquired in circular SAR mode through 8 passes at different altitudes, with elevation angles of [45.66°, 44.01°, 43.92°, 44.18°, 44.14°, 43.53°, 43.01°, 43.06°]. The combination of circular observation aperture and altitude diversity enabled 3D imaging of the entire scene. We select data from two SUVs and two sedans. Figure 17 illustrates the scene, radar flight passes, and vehicle photos.
@@ III-C1 Real-world dataset
Figure 18: Imaging results of different methods using real-world data. Imaging using 12 sub-aspects
Figure 19: Imaging results of different methods using real-world data. Imaging using 10 sub-aspects
Figure 20: Imaging results of different methods using real-world data. Imaging using 8 sub-aspects
Figure 21: Imaging results of different methods using real-world data. Imaging using 6 sub-aspects
Figure 22: Imaging results of different methods using real-world data. Imaging using 4 sub-aspects
III-C2 Imaging results
Figure 18 to 22 present the imaging results on real-world data with varying number of sub-aspects. To facilitate observation, the visualization angle is set to (0°, 0°) when presenting the results. The BP method produces results heavily affected by noise and side lobes, which can hardly be recognized. The CS method alleviates these issues, but the reconstructions still appear as chaotic point clouds, leaving vehicle features unrecognizable, especially when using highly sparse data. For example, when the number of sub-aspects is 4, traditional methods completely fail, producing totally disordered point clusters without any discernible or meaningful vehicle features.
In contrast, radar image-supervised DL methods can reconstruct vaguely discernible vehicle contours. However, as the number of aspects decreases, these contours deteriorates. For example, when the number of sub-aspects is reduced to 4, SACNet and UNet3D show significant omissions in reconstructing the roof of Jeep. After integrating cross-modal supervision, SACNet and UNet3D exhibit weak generalization when processing real-world data, struggling to accurately reconstruct the vehicle’s geometry. Particularly, when the number of sub-aspects is 4, SACNet fails completely and cannot output any visible features.
Tables IV-V and Figure 23 show the evaluation metrics. Compared to traditional methods and DL methods supervised with radar images, CMAR-Net achieved improvements of $>$100% in IoU, 89.36% in CE, 75.83% in PSNR and 47.85% in SSIM. When compared to DL methods with cross-modal supervision, CMAR-Net shows enhancements of 86.42% in IoU, 67.99% in CE, 38.81% in PSNR and 10.75% in SSIM.
Figure 23: Image quality metrics-number of sub-aspect curves for different methods on real-world data.
@@ III-C2 Imaging results
Methods Number of sub-aspect
@@
Methods Number of sub-aspect
@@
III-C3 Ablation study
The 7th and 8th sets of data in the Figure 18-22 and table IV-V show the imaging results and evaluation metrics after ablating the PRP unit and data augmentation strategy. Without the PRP unit, the vehicle images become blurry and suffer greater pixel loss as the number of aspects decreases. And in the absence of data augmentation, the network consistently reconstructs all vehicle type as sedan type, revealing severe overfitting. The PRP unit design improves SSIM, PSNR, IoU, and CE by 8.25%, 21.25%, 24.15%, and 42.82%, respectively. Data augmentation contributes improvements of 8.65%, 19.92%, 42.94%, and 46.75%. These results clearly demonstrate the effectiveness of both the PRP unit and data augmentation strategy.
IV Conclusion
In this paper, we propose a DL-based multi-baseline SAR 3D reconstruction method that utilizes only sparse aspect data. We introduce the concept of cross-modal learning and design a DNN tailored for 3D SAR reconstruction—CMAR-Net. By leveraging differentiable rendering, CMAR-Net can be effectively trained, addressing the key challenge of supervising 3D reconstruction network with 2D optical images. Extensive experiments on both simulated and real-world datasets show that CMAR-Net significantly outperforms all existing baseline methods, including traditional imaging algorithms and DL-based approaches, in terms of image quality and generalization. Moreover, the existing DL-based SAR reconstruction methods, adopting our proposed cross-modal learning paradigm, experience a significant performance improvement, which highlights the superiority and versatility. Last but not least, our method eliminates the need for time-consuming preprocessing of full-aperture data. Instead, it only needs 2D images obtained through computer rendering or camera capture. This significantly reduces the difficulty of dataset construction, greatly enhancing the practicality of 3D SAR reconstruction technology.
Our approach offers new insights for 3D SAR researchers and is expected to comprehensively improve the research landscape of 3D SAR imaging. The technique has the potential to be applied to fields such as aerial remote sensing, terrain mapping, and obstacle detection. In the future, researchers can expand the scope to 3D reconstruction of other scenes (e.g., buildings, streets), or explore fusion with other heterogeneous data (e.g., infrared data), and so on.